story problem "yuck"
The cost of a computer decreases from $2500 to $1200 after 3 months. find the percent decrease.
I got 0.2% but that is not correct.
Please help.

Answer by fractalier(6550)

Percentage decrease = (new - old) / old times 100% so we get
% Decr = (1200 - 2500) / 2500 x 100%
% Decr = -1300 / 2500 x 100%
% Decrease = -52%
or a 52% decrease in cost